construct a 2 by 2 table, calculate predictive value positive and predictive value negative of the EIA in a hypothetical population of 1,000,000 blood donors. using a separate 2 by 2 table calculate PVP and PVN for a population of 1000 ill patients. Assume that the actual prevalence of S. heidelberg among blood donors is 0.04% (0.0004 ) and that of people who ate foster's chicken is 10.0% (0.10).
Answer #1

Wednesday (100-256) Week 16 able April 19 Wocreado 2by2 Table April Test Result T Present Absent Total Positive 380 19992 203

PVP = 95/113 = 84.1% PVN = 882/887 99.4%